There are 4 Republicans and 5 Democrats. How many committees of 4 members can be formed if there must be 2 Republicans and 2 Democrats in each committee but a particular Republican must be in the committee?


1
Expert's answer
2021-11-15T17:56:43-0500

Number of ways to choose 1 Republican from 3 and 2 Democrat from 5

= (3C1)*(5C2) = 3*10 = 30

(as one Republican already selected , therefore we need to select 1 more from remaining 3)

Ans : 30 committee